# Downlink and Uplink ISAC in Continuous-Aperture Array (CAPA) Systems

### v1

@@ VI Conclusion
This article proposed an analytically tractable framework for both downlink and uplink CAPA-based ISAC by leveraging continuous operators to model the continuous nature of the EM field. Based on this framework, we analyzed the sensing and communication performance in terms of the SR and CR. For downlink ISAC, we proposed three continuous beamforming designs: the S-C design, the C-C design, and the Pareto-optimal design. For uplink ISAC, we investigated the S-C SIC and C-C SIC methods. By employing a signal subspace-based approach, we derived closed-form expressions for optimal continuous beamformers as well as the achievable CRs, SRs, and rate regions in both downlink and uplink scenarios. Numerical results demonstrated that CAPA-based ISAC achieves higher CRs and SRs than both traditional SPDA-based ISAC and CAPA-based FDSAC systems. These findings validate the significant advantages of CAPA, making it a promising technology for next-generation ISAC systems.
@@
$\begin{split}h_{\rm{d}}({\mathbf{t}})&=\frac{\mathrm{j}\eta k_{0}\sqrt{\frac{r_{\mathrm{c}}\Psi_{\mathrm{c}}}{4\pi}}\mathrm{e}^{-\mathrm{j}k_{0}(x^{2}+z^{2}-2r_{\mathrm{c}}\left(\Phi_{\mathrm{c}}x+\Theta_{\mathrm{c}}z\right)+r_{\mathrm{c}}^{2})^{\frac{1}{2}}}}{(x^{2}+z^{2}-2r_{\mathrm{c}}\left(\Phi_{\mathrm{c}}x+\Theta_{\mathrm{c}}z\right)+r_{\mathrm{c}}^{2})^{\frac{3}{4}}}\\ &=\hat{h}_{\mathrm{c}}(x,z),\ x\in[0,L_{x}],z\in[-{L_{z}}/{2},{L_{z}}/{2}].\end{split}$ (A2)
@@
$\begin{split}&\int_{\mathcal{A}_{\mathrm{t}}}{\left|h_{\rm{d}}({\mathbf{t}})\right|^{2}}\mathrm{d}\mathbf{t}=\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}{\int_{0}^{L_{x}}{\hat{h}_{\mathrm{c}}(x,z)\hat{h}_{\mathrm{c}}^{*}(x,z)\mathrm{d}x\mathrm{d}z}}\\ &=\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}\int_{0}^{L_{x}}\frac{\frac{1}{4\pi}\eta^{2}k_{0}^{2}r_{\mathrm{c}}\Psi_{\mathrm{c}}\mathrm{d}x\mathrm{d}z}{\left(x^{2}\!+\!z^{2}\!-\!2r_{\mathrm{c}}\!\left(\Phi_{\mathrm{c}}x\!+\!\Theta_{\mathrm{c}}z\right)\!+\!r_{\mathrm{c}}^{2}\right)^{\frac{3}{2}}}.\end{split}$ (A3)
The inner integral can be evaluated using [45, Eq. (2.264.5)], and the outer integral using[45, Eq. (2.284)], which yields $\int_{\mathcal{A}_{\mathrm{t}}}{\left|h_{\rm{d}}({\mathbf{t}})\right|^{2}}\mathrm{d}\mathbf{t}=g_{\mathrm{d}}$. Thus, the results of (39) follow directly from this calculation.
-B Proof of Lemma 1
@@
-D Proof of Lemma 3
Inserting ${\mathbf{v}}=\mathbf{v}_{\star}$ into the objective function of problem (71), we obtain
$\max_{v_{\rm{s}}({\mathbf{r}})}\frac{\lvert\int_{{\mathcal{A}}_{\rm{r}}}v_{\rm{s}}^{*}({\mathbf{r}})a_{\rm{r}}(\mathbf{r}){\rm{d}}{\mathbf{r}}\rvert^{2}}{\sigma_{{\rm{c}},\ell}^{2}}$ $=\frac{1}{\sigma^{2}}\frac{{\mathbf{v}}_{\star}^{\mathsf{H}}{\mathbf{a}_{\rm{r}}}{\mathbf{a}_{\rm{r}}^{\mathsf{H}}}{\mathbf{v}}_{\star}}{{\mathbf{v}}_{\star}^{\mathsf{H}}(\tilde{\gamma}_{\mathrm{c}}{\mathbf{h}}_{\rm{u}}{\mathbf{h}}_{\rm{u}}^{\mathsf{H}}+{\mathbf{I}_{2}}){\mathbf{v}}_{\star}}$ (A19)
$=\frac{{\mathbf{a}}_{\rm{r}}^{\mathsf{H}}(\tilde{\gamma}_{\mathrm{c}}{\mathbf{h}}_{\rm{u}}{\mathbf{h}}_{\rm{u}}^{*}+{\mathbf{I}}_{2})^{-1}{\mathbf{a}_{\rm{r}}}}{\sigma^{2}}.$ (A20)
Using the Woodbury matrix identity, we can rewrite (A20) as follows:
[…]
$=\frac{1}{\sigma^{2}}\left(\lVert{\mathbf{a}}_{\rm{r}}\rVert^{2}-\frac{\tilde{\gamma}_{\mathrm{c}}\lvert{\mathbf{h}}_{\rm{u}}^{\mathsf{H}}{\mathbf{a}}_{\rm{r}}\rvert^{2}}{1+\tilde{\gamma}_{\mathrm{c}}\lVert{\mathbf{h}_{\rm{u}}}\rVert^{2}}\right).$ (A22)
Furthermore, we have $\|\mathbf{h}_{\mathrm{u}}\|^{2}=\int_{\mathcal{A}_{\mathrm{r}}}|h_{\mathrm{u}}(\mathbf{r})|^{2}\mathrm{d}\mathbf{r}$, $\|\mathbf{a}_{\mathrm{r}}\|^{2}=\int_{\mathcal{A}_{\mathrm{r}}}|a_{\mathrm{r}}(\mathbf{r})|^{2}\mathrm{d}\mathbf{r}$, and $\mathbf{h}_{\mathrm{u}}^{\mathsf{H}}\mathbf{a}_{\mathrm{r}}=\int_{\mathcal{A}_{\mathrm{r}}}{a_{\mathrm{r}}(\mathbf{r})h_{\mathrm{u}}^{*}(\mathbf{r})}\mathrm{d}\mathbf{r}$. Thus, the results in (73) follow immediately.

### v2

@@ VI Conclusion
This article proposed an analytically tractable framework for both downlink and uplink CAPA-based ISAC by leveraging continuous operators to model the continuous nature of the EM field. Based on this framework, we analyzed the sensing and communication performance in terms of the SR and CR. For downlink ISAC, we proposed three continuous beamforming designs: the S-C design, the C-C design, and the Pareto-optimal design. For uplink ISAC, we investigated the S-C SIC and C-C SIC methods. By employing a low-complexity signal subspace-based approach, we derived closed-form expressions for optimal continuous beamformers as well as the achievable CRs, SRs, and rate regions in both downlink and uplink scenarios. Numerical results demonstrated that CAPA-based ISAC achieves better sensing and communication performance than both traditional SPDA-based ISAC and CAPA-based FDSAC systems. These findings validate the significant advantages of CAPA, making it a promising technology for next-generation ISAC systems. As an extension, future work will consider more general CAPA-based ISAC scenarios involving multiple CUs and targets, where joint beamforming and interference management become critical challenges.
@@
By defining $\hat{\mathbf{h}}_{\mathrm{s}}\triangleq\sqrt{P}\hat{v}_{\mathrm{s}}\hat{w}_{\mathrm{s}}\mathbf{s}$, (29) is rewritten as
$\hat{\mathbf{y}}_{\mathrm{s}}=\hat{\mathbf{h}}_{\mathrm{s}}\beta_{\mathrm{s}}+\hat{\mathbf{n}}_{\mathrm{s}}.$ (A1)
Given the observation $\hat{\mathbf{y}}_{\mathrm{s}}$, the MSE of the estimator $f_{\mathrm{est}}\left(\hat{\mathbf{y}}_{\mathrm{s}}\right)$ for the reflection coefficient $\beta_{\mathrm{s}}$ is expressed as follows:
$\mathsf{MSE}=\mathbbmss{E}\left\{\left|f_{\mathrm{est}}\left(\hat{\mathbf{y}}_{\mathrm{s}}\right)-\beta_{\mathrm{s}}\right|^{2}\right\}.$ (A2)
It is well known that the minimum value of (A2) is attained by the conditional mean estimator $f_{\mathrm{est}}\left(\hat{\mathbf{y}}_{\mathrm{s}}\right)=\mathbbmss{E}\left\{\beta_{\mathrm{s}}|\hat{\mathbf{y}}_{\mathrm{s}}\right\}$. To calculate the conditional mean, we need to obtain the conditional probability density function (PDF) of $\beta_{\mathrm{s}}$ given $\hat{\mathbf{y}}_{\mathrm{s}}$, denoted as $f_{\beta_{\mathrm{s}}|\hat{\mathbf{y}}_{\mathrm{s}}}\left(x|\mathbf{y}\right)$. Based on Bayes’ theorem, we have
$f_{\beta_{\mathrm{s}}|\hat{\mathbf{y}}_{\mathrm{s}}}\left(x|\mathbf{y}\right)=\frac{f_{\beta_{\mathrm{s}}}\left(x\right)f_{\hat{\mathbf{y}}_{\mathrm{s}}|\beta_{\mathrm{s}}}\left(\mathbf{y}|x\right)}{f_{\hat{\mathbf{y}}_{\mathrm{s}}}\left(\mathbf{y}\right)},$ (A3)
where $f_{\beta_{\mathrm{s}}}\left(\cdot\right)$, $f_{\hat{\mathbf{y}}_{\mathrm{s}}|\beta_{\mathrm{s}}}\left(\cdot|\cdot\right)$, and $f_{\hat{\mathbf{y}}_{\mathrm{s}}}\left(\cdot\right)$ denote the PDFs of $\beta_{\mathrm{s}}$, $\hat{\mathbf{y}}_{\mathrm{s}}$ conditioned on $\beta_{\mathrm{s}}$, and $\hat{\mathbf{y}}_{\mathrm{s}}$, respectively. Given that $\beta_{\mathrm{s}}\sim\mathcal{C}\mathcal{N}(0,\alpha_{\mathrm{s}})$ and $\hat{\mathbf{n}}_{\mathrm{s}}\sim\mathcal{C}\mathcal{N}(0,\hat{\sigma}_{\mathrm{s}}^{2})$ with $\hat{\sigma}_{\mathrm{s}}^{2}\triangleq\sigma_{\mathrm{s}}^{2}\int_{\mathcal{A}_{\mathrm{r}}}{\left|v_{\mathrm{s}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}$, we have
$f_{\beta_{\mathrm{s}}}\left(x\right)=\frac{1}{\pi\alpha_{\mathrm{s}}}\mathrm{e}^{-\frac{\left|x\right|^{2}}{\alpha_{\mathrm{s}}}},$ (A4a)
$f_{\hat{\mathbf{y}}_{\mathrm{s}}|\beta_{\mathrm{s}}}(\mathbf{y}|x)=\frac{1}{\left(\pi\hat{\sigma}_{\mathrm{s}}^{2}\right)^{L}}\exp\Big{(}\!-{\hat{\sigma}_{\mathrm{s}}^{-2}}{\big{(}\mathbf{y}-\hat{\mathbf{h}}_{\mathrm{s}}x\big{)}^{\mathsf{H}}\big{(}\mathbf{y}-\hat{\mathbf{h}}_{\mathrm{s}}x\big{)}}\Big{)},$ (A4b)
$f_{\hat{\mathbf{y}}_{\mathrm{s}}}\left(\mathbf{y}\right)=\frac{\exp\left(-\mathbf{y}^{\mathsf{H}}\big{(}\alpha_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}^{\mathsf{H}}+\hat{\sigma}_{\mathrm{s}}^{2}\mathbf{I}_{L}\big{)}^{-1}\mathbf{y}\right)}{\pi^{L}\det\big{(}\alpha_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}^{\mathsf{H}}+\hat{\sigma}_{\mathrm{s}}^{2}\mathbf{I}_{L}\big{)}}$
=$~{}~{}~{}~{}~{}\overset{(a)}{=}\frac{\exp\left(-\hat{\sigma}_{\mathrm{s}}^{-2}\mathbf{y}^{\mathsf{H}}\left(\mathbf{I}_{L}-\frac{\alpha_{\mathrm{s}}}{\hat{\sigma}_{\mathrm{s}}^{2}+\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}}\hat{\mathbf{h}}_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}^{\mathsf{H}}\right)\mathbf{y}\right)}{\left(\pi\hat{\sigma}_{\mathrm{s}}^{2}\right)^{L}\big{(}1+{\hat{\sigma}_{\mathrm{s}}^{-2}}{\alpha_{\mathrm{s}}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}\big{)}},$ (A4c)
where equality $(a)$ follows from the matrix determinant lemma, i.e., $\det\big{(}\alpha_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}^{\mathsf{H}}+\hat{\sigma}_{\mathrm{s}}^{2}\mathbf{I}_{L}\big{)}=\left(\hat{\sigma}_{\mathrm{s}}^{2}\right)^{L}\big{(}1+{\hat{\sigma}_{\mathrm{s}}^{-2}}{\alpha_{\mathrm{s}}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}\big{)}$ and Woodbury matrix identity, i.e., $\big{(}\alpha_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}^{\mathsf{H}}+\hat{\sigma}_{\mathrm{s}}^{2}\mathbf{I}_{L}\big{)}^{-1}=\hat{\sigma}_{\mathrm{s}}^{-2}\big{(}\mathbf{I}_{L}-\frac{\alpha_{\mathrm{s}}}{\hat{\sigma}_{\mathrm{s}}^{2}+\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}}\hat{\mathbf{h}}_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}^{\mathsf{H}}\big{)}$. Substituting (A4) into (A3) gives
$\begin{split}&f_{\beta_{\mathrm{s}}|\hat{\mathbf{y}}_{\mathrm{s}}}\left(x|\mathbf{y}\right)=\frac{\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}+\hat{\sigma}_{\mathrm{s}}^{2}}{\pi\alpha_{\mathrm{s}}\hat{\sigma}_{\mathrm{s}}^{2}}\times\\ &~{}~{}\exp\!\left(\!-\frac{\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}+\hat{\sigma}_{\mathrm{s}}^{2}}{\alpha_{\mathrm{s}}\hat{\sigma}_{\mathrm{s}}^{2}}\left|x-\frac{\alpha_{\mathrm{s}}}{\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}+\hat{\sigma}_{\mathrm{s}}^{2}}\hat{\mathbf{h}}_{\mathrm{s}}^{\mathsf{H}}\mathbf{y}\right|^{2}\right).\end{split}$ (A5)
As a result, we can obtain the conditional mean estimator as
$f_{\mathrm{est}}\left(\hat{\mathbf{y}}_{\mathrm{s}}\right)=\mathbbmss{E}\left\{\beta_{\mathrm{s}}|\hat{\mathbf{y}}_{\mathrm{s}}\right\}$ $=\int_{0}^{\infty}{xf_{\beta_{\mathrm{s}}|\hat{\mathbf{y}}_{\mathrm{s}}}\left(x|\mathbf{y}\right)\mathrm{d}}x$
$=\frac{\alpha_{\mathrm{s}}}{\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}+\hat{\sigma}_{\mathrm{s}}^{2}}\hat{\mathbf{h}}_{\mathrm{s}}^{\mathsf{H}}\mathbf{y}.$ (A6)
Inserting (A1) and (-A) into (A2) yields
$\mathsf{MSE}$ $=\mathbbmss{E}\left\{\left|\frac{\alpha_{\mathrm{s}}}{\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}+\hat{\sigma}_{\mathrm{s}}^{2}}\hat{\mathbf{h}}_{\mathrm{s}}^{\mathsf{H}}\left(\hat{\mathbf{h}}_{\mathrm{s}}\beta_{\mathrm{s}}+\hat{\mathbf{n}}_{\mathrm{s}}\right)-\beta_{\mathrm{s}}\right|^{2}\right\}$
$=\mathbbmss{E}\left\{\left|\frac{-\sigma_{\mathrm{s}}^{2}\beta_{\mathrm{s}}}{\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}+\hat{\sigma}_{\mathrm{s}}^{2}}+\frac{\alpha_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}^{\mathsf{H}}\hat{\mathbf{n}}_{\mathrm{s}}}{\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}+\hat{\sigma}_{\mathrm{s}}^{2}}\right|^{2}\right\}.$ (A7)
Since $\beta_{\mathrm{s}}$ and $\hat{\mathbf{n}}_{\mathrm{s}}$ are mutually independent, (-A) can be further simplified as follows:
$\mathsf{MSE}$ $=\mathbbmss{E}\left\{\left|\frac{-\sigma_{\mathrm{s}}^{2}\beta_{\mathrm{s}}}{\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}+\hat{\sigma}_{\mathrm{s}}^{2}}\right|^{2}\right\}+\mathbbmss{E}\left\{\left|\frac{\alpha_{\mathrm{s}}\hat{\mathbf{h}}_{\mathrm{s}}^{\mathsf{H}}\hat{\mathbf{n}}_{\mathrm{s}}}{\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}+\hat{\sigma}_{\mathrm{s}}^{2}}\right|^{2}\right\}$
$=\frac{\alpha_{\mathrm{s}}}{1+\alpha_{\mathrm{s}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}\hat{\sigma}_{\mathrm{s}}^{-2}}=\frac{\alpha_{\mathrm{s}}}{1+L\overline{\gamma}_{\mathrm{s}}\frac{\left|\hat{v}_{\mathrm{s}}\right|^{2}\left|\hat{w}_{\mathrm{s}}\right|^{2}}{\int_{\mathcal{A}_{\mathrm{r}}}{\left|v_{\mathrm{s}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}}}.$ (A8)
By observing (-A) and (33), the optimal sensing detector that minimizes the MSE follows ${{v}}_{\rm{s}}({\mathbf{r}})\propto a_{\rm{r}}(\mathbf{r})$, which corresponds to the detector that maximizes the SR. Consequently, the minimum MSE (MMSE) can be expressed as follows:
$\mathsf{MMSE}({{w}}({\mathbf{t}}))=\frac{\alpha_{\mathrm{s}}}{1+L\overline{\gamma}_{\mathrm{s}}\left|\hat{w}_{\mathrm{s}}\right|^{2}\int_{\mathcal{A}_{\mathrm{r}}}{\left|a_{\mathrm{r}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}}.$ (A9)
The BCRB provides a fundamental lower bound on the MSE achievable by any unbiased Bayesian estimator. Unlike the conventional CRB which assumes deterministic unknowns, the BCRB incorporates prior information about the parameter being estimated. Based on [59], the Bayesian Fisher information for estimating $\beta_{\mathrm{s}}$ is given by
${\mathsf{BFI}}={\mathbbmss{E}}\left\{\left|\frac{\partial\ln{f_{\beta_{\mathrm{s}}}\left(\beta_{\mathrm{s}}\right)}}{\partial\beta_{\mathrm{s}}}\right|^{2}\right\}+{\mathbbmss{E}}\left\{\left|\frac{\partial\ln{f_{\hat{\mathbf{y}}_{\mathrm{s}}|\beta_{\mathrm{s}}}(\hat{\mathbf{y}}_{\mathrm{s}}|\beta_{\mathrm{s}})}}{\partial\beta_{\mathrm{s}}}\right|^{2}\right\}.$ (A10)
Referring to (A4) and performing some basic mathematical manipulations, we obtain
${\mathbbmss{E}}\left\{\left|\frac{\partial\ln{f_{\beta_{\mathrm{s}}}\left(\beta_{\mathrm{s}}\right)}}{\partial\beta_{\mathrm{s}}}\right|^{2}\right\}=\frac{1}{\alpha_{\mathrm{s}}},$ (A11a)
${\mathbbmss{E}}\left\{\left|\frac{\partial\ln{f_{\hat{\mathbf{y}}_{\mathrm{s}}|\beta_{\mathrm{s}}}(\hat{\mathbf{y}}_{\mathrm{s}}|\beta_{\mathrm{s}})}}{\partial\beta_{\mathrm{s}}}\right|^{2}\right\}=\frac{1}{\hat{\sigma}_{\mathrm{s}}^{2}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}.$ (A11b)
It follows that the BCRB can be written as follows [59]:
$\mathsf{BCRB}=\frac{1}{{\mathsf{BFI}}}=\frac{1}{\frac{1}{\alpha_{\mathrm{s}}}+\frac{1}{\hat{\sigma}_{\mathrm{s}}^{2}}\lVert\hat{\mathbf{h}}_{\mathrm{s}}\rVert^{2}}=\mathsf{MSE},$ (A12)
where the last equality is due to (-A). By further referring to (A9), we find that the minimum BCRB is given by $\mathsf{BCRB}({{w}}({\mathbf{t}}))=\frac{\alpha_{\mathrm{s}}}{1+L\overline{\gamma}_{\mathrm{s}}\left|\hat{w}_{\mathrm{s}}\right|^{2}\int_{\mathcal{A}_{\mathrm{r}}}{\left|a_{\mathrm{r}}(\mathbf{r})\right|^{2}}\mathrm{d}\mathbf{r}}$. This implies that, under the considered model, the MMSE estimator attains the BCRB, making it an efficient estimator.
By comparing the results of (A9) and (A12) with the definition of the SR given in (35), it is clearly shown that both the MSE and BCRB are negatively correlated to the SR, which means that maximizing SR is equivalent to minimizing the MMSE and BCRB. This completes the proof.
-B Proof of Theorem 2
@@
$\begin{split}h_{\rm{d}}({\mathbf{t}})&=\frac{\mathrm{j}\eta k_{0}\sqrt{\frac{r_{\mathrm{c}}\Psi_{\mathrm{c}}}{4\pi}}\mathrm{e}^{-\mathrm{j}k_{0}(x^{2}+z^{2}-2r_{\mathrm{c}}\left(\Phi_{\mathrm{c}}x+\Theta_{\mathrm{c}}z\right)+r_{\mathrm{c}}^{2})^{\frac{1}{2}}}}{(x^{2}+z^{2}-2r_{\mathrm{c}}\left(\Phi_{\mathrm{c}}x+\Theta_{\mathrm{c}}z\right)+r_{\mathrm{c}}^{2})^{\frac{3}{4}}}\\ &=\tilde{h}_{\mathrm{c}}(x,z),\ x\in[0,L_{x}],z\in[-{L_{z}}/{2},{L_{z}}/{2}].\end{split}$ (A14)
@@
$\begin{split}&\int_{\mathcal{A}_{\mathrm{t}}}{\left|h_{\rm{d}}({\mathbf{t}})\right|^{2}}\mathrm{d}\mathbf{t}=\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}{\int_{0}^{L_{x}}{\tilde{h}_{\mathrm{c}}(x,z)\tilde{h}_{\mathrm{c}}^{*}(x,z)\mathrm{d}x\mathrm{d}z}}\\ &=\int_{-\frac{L_{z}}{2}}^{\frac{L_{z}}{2}}\int_{0}^{L_{x}}\frac{\frac{1}{4\pi}\eta^{2}k_{0}^{2}r_{\mathrm{c}}\Psi_{\mathrm{c}}\mathrm{d}x\mathrm{d}z}{\left(x^{2}\!+\!z^{2}\!-\!2r_{\mathrm{c}}\!\left(\Phi_{\mathrm{c}}x\!+\!\Theta_{\mathrm{c}}z\right)\!+\!r_{\mathrm{c}}^{2}\right)^{\frac{3}{2}}}.\end{split}$ (A15)
The inner integral can be evaluated using [60, Eq. (2.264.5)], and the outer integral using[60, Eq. (2.284)], which yields $\int_{\mathcal{A}_{\mathrm{t}}}{\left|h_{\rm{d}}({\mathbf{t}})\right|^{2}}\mathrm{d}\mathbf{t}=g_{\mathrm{d}}$. Thus, the results of (39) follow directly from this calculation.
-C Proof of Lemma 1